You have been asked to develop a pro forma statement of cash flow for an office plaza. The information given to you is listed below:

Property Information:

Age 8 years

Rentable Space 300,000 sq.ft.

# Stories 15

# Tenants 40

Financial Information:

Base Rent Avg. :$20 per sq.ft.

Other Income/Parking: $1.50 per sq.ft.

Expenses Recoverable from Tenants: $2.50 per sq.ft.

Current Vacancy: equals to 5% of base rents

Operating Expenses $4,225,100

What is the net operating income for the office plaza based on the pro forma statement of cash flow for a base year you developed?
Answer #1
RECEIPTS:
Base rent = 300000*20*95% = $    57,00,000
Other income/Parking = 300000*95%*1.5 = $      4,27,500
Recoveries towards expenses = 300000*95%*2.5 = $      7,12,500
Total receipts $ 68,40,000
PAYMENTS:
Operating expenses $ 42,25,100
Net operating income $ 26,14,900